OpenStudy (anonymous):

Arrange the numbers from least to greatest. Explainn you'r method. -2 1/3 -26/11 -2.360 360 360 -2 2/5 -2.36

OpenStudy (cwrw238):

convert the fractions to decimals - then the order will be obvious

OpenStudy (anonymous):

yeah so, idk how to do that :"(

OpenStudy (cwrw238):

1/3 : divide 1 by 3 gives 0.3333 26/11 - division gives 2 . 3636 2/5 = 0.4

OpenStudy (anonymous):

its - 2 1/3 ;s

OpenStudy (cwrw238):

s0 -2 1.3 = -2.3333 -2 2/5 = -2.4 -26/11 = -2.3636

OpenStudy (cwrw238):

so least number is -2.4 then -2.3636 -2.36 -2.360 - can you continue from here?

OpenStudy (cwrw238):

-2.36 and -2.360 are really the same

OpenStudy (anonymous):

wheres -2.4 coming from and -3.3636 :S:S

OpenStudy (cwrw238):

I explained above -2 2/5 = -2.4 because 2/5 = 0.4 - check my earlier post

OpenStudy (anonymous):

so that ones the least ?

OpenStudy (anonymous):

isnt - 2 1/3 less than -26/11